OpenStudy (australopithecus):

Find the equation of the tangent plane of the surface z^2 + x^2 − 4xy + y^2 = 2 at the point (1, 1, 2).

OpenStudy (australopithecus):

I know the equation is normally \[z - z_0 = f_x(x_0,y_0)(x-x_0) - f_y(x_0,y_0)(y-x_0)\] would I just use implicit differentiation to solve? so take the derivative in respect to z?

OpenStudy (anonymous):

according to the equation u gave us we just need to computing \(f_x(x_0,y_0)\) and \(f_y(x_0,y_0)\) \(z=f(x,y) \) for \(f_x(x_0,y_0)\) take partial derivative with respect to \(x\) and for \(f_y(x_0,y_0)\) take partial derivative wrt y

OpenStudy (anonymous):

and yes implicit differentiation...

OpenStudy (australopithecus):

so I got \[f_x(x,y) = \frac{-4x+2y}{-2z}\] and \[f_y(x,y) = \frac{2x-4y}{-2z}\]

OpenStudy (australopithecus):

I'm pretty sure i made a mistake

OpenStudy (anonymous):

yep

OpenStudy (australopithecus):

as I got z = (1/2)(x+y) +1 which seems incorrect

OpenStudy (anonymous):

see when u take derivative with respect to x u must treat y like a constant

OpenStudy (anonymous):

now u tell me what is \(z'_y\)

OpenStudy (australopithecus):

\[z_y = \frac{-4x + 2y}{-2z}\] I got the same thing as last time

OpenStudy (australopithecus):

so (2x-y)/z

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

i made a little mistake too....lol \[z'_x=\frac{2y-x}{z} \]

OpenStudy (australopithecus):

this is in the list of possible answers A. 2z − x − y = 2 B. 2z + x + y = 6 C. z + x + y = 4 D. z = 2 E. 2x + 2y − z = 2 F. z − x − y = 0

OpenStudy (anonymous):

o.O i got 2z-x+y=4 what about u?

OpenStudy (australopithecus):

I got x + y + 2= 2z

OpenStudy (australopithecus):

ugh my brain is so dead right now

OpenStudy (australopithecus):

2z - x - y = 2

OpenStudy (anonymous):

we're both wrong.....!!!

OpenStudy (australopithecus):

zx(x,y) = 1/2 zy(x,y) = 1/2 (1/2)(x-1) + (1/2)(y-1) + 2 = z

OpenStudy (australopithecus):

(1/2)(x-1) + (1/2)(y-1) + 2 = z = (1/2)x - 1/2 + (1/2)y - 1/2 + 2 = z = z - 1 =1/2(x+y) = 2z - 2 = x+y = 2z - x - y = 2

OpenStudy (anonymous):

oh .... there is a little typo in your first reply z-z0=.....+.....(y-y0) not ........-.......(y-x0)

OpenStudy (australopithecus):

lol sorry about that I messed up the formula

OpenStudy (australopithecus):

none the less they should be the same x_0 = 1, y_0 = 1

OpenStudy (australopithecus):

so that formula works for this question

OpenStudy (anonymous):

so answer is 2z-x-xy=2

OpenStudy (australopithecus):

so it seems thanks for the help

OpenStudy (anonymous):

yw :)

OpenStudy (australopithecus):

Man I need sleep